P(x)=x+1/6 where x=0,1,2.


1. P(x=2)

2. P(x>1)

3. P(<1)

Solution







1
Expert's answer
2021-09-27T23:29:36-0400
"P(X=2)=\dfrac{2+1}{6}=\dfrac{1}{2}"

"P(X>1)=P(X=2)=\dfrac{2+1}{6}=\dfrac{1}{2}"

"P(X<1)=P(X=0)=\dfrac{0+1}{6}=\dfrac{1}{6}"